Two equal and opposite forces each of magnitude 40N are applied to the rim of a wheel makes 12 revolutions while the torque is kept constant. Calculate the angle of revolution, the torque and the workdone


1
Expert's answer
2021-04-09T06:38:17-0400
"\theta=12(2\pi)=24\pi \ rad\\\tau=2Tr=2(40)(2.5)=200\ Nm\\
W=\tau \theta=(24\pi)(200)=15080\ J"